OpenStudy (anonymous):

the formula for compound intrest is given by A=P(1+r)^t where A is the total amount of money, P is the principalinvested, r is the intrest rate, and t is the time in years. If $5000 invested at 6.5% intrest compounded annually yields $25,000, for how many years was the money invested?

OpenStudy (jamesj):

Given this equation A=P(1+r)^t how do you find t in terms of the other variables?

OpenStudy (jamesj):

Well, A/P = (1+r)^t hence taking the natural log of both sides ln(A/P) = ln( (1+r)^t ) = t . ln(1+r) therefore t = ln(A/P) / ln(1+r)

OpenStudy (jamesj):

Now substitute all of your values of A, P and r and find t.

OpenStudy (anonymous):

can you explain it

OpenStudy (jamesj):

which step of the algebra don't you understand: step 1: A/P = (1+r)^t hence taking the natural log of both sides step 2: ln(A/P) = ln( (1+r)^t ) step 3: = t . ln(1+r) therefore step 4: t = ln(A/P) / ln(1+r)

OpenStudy (anonymous):

what does r equal

OpenStudy (jamesj):

r is the interest rate 6.5%. As ever, you need to express that in decimal format. Hence r = 0.065

OpenStudy (anonymous):

so would i substitue like this. ln(5000)/ln(1+1.065)

OpenStudy (jamesj):

Almost ln(A/P), not ln(P)

OpenStudy (anonymous):

so ln(25,000/5000)/ln(1+0.065)

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

is the answer 4.69